# CoSteer: Collaborative Decoding-Time Personalization via Local Delta Steering

### v2

@@ A.3 Proof
Appendix B Experiment Detail
@@ B.1 Hyperparameters and Settings
To ensure the results are easily reproducible, we set do_sample=False for all methods. For Qwen3 series, we set enable_thinking=False. Hyperparameters of Costeer are listed in Table 4
@@ B.2 Detailed Results
Qwen 8B-0.6B
SLM w/o 6.84 6.64 37.13 20.41 21.00 24.41 12.83 12.39 24.19 11.88 15.74 .5710 .5359 .5981 .5937
Qwen 32B-0.6B
SLM w/o 6.84 6.64 37.13 20.41 21.00 24.41 12.83 12.39 24.19 11.88 15.74 .5710 .5359 .5981 .5937
CoSteer 8.52 8.56 43.74 23.69 30.97 30.89 13.14 18.87 26.75 13.17 21.32 .7858 .7883 .6873 .7790
Table 6: Comparative performance across eight personalized content generation and preference alignment tasks. Metrics include overall (Ovl) and personalized (Per) scores for Cogenesis, ROUGE-1/-L (R-1/-L) and METEOR (MET) for Longlamp datasets, and averaged alignment scores for four user preferences. Bold entries indicate that CoSteer outperforms the three baseline methods compared against. Gray values represent the privacy-violating near-upper-bound performance, yet underlined CoSteer values surpasses these incompatible references.
@@ B.3 Statistical Significance Testing
Qwen 7B-1.5B
SLM w/o 6.63 6.21 36.48 17.74 27.57 20.40 10.39 10.39 25.21 11.09 17.51 .5441 .5154 .6104 .5632
CoSteer 8.44* 8.50* 42.98* 23.61* 28.20* 32.72* 15.92* 20.36* 25.93 12.38* 22.84* .7915* .7608* .7404* .7885*
Qwen 32B-7B
SLM w/o 8.00 7.63 39.81 20.53 25.56 30.15 14.04 17.71 27.64 11.93 21.49 .7058 .6931 .7039 .7183
CoSteer 8.78* 8.64 45.41* 26.04* 33.52* 34.88* 15.89* 26.51* 30.10* 14.52* 24.20* .8589 .8532* .7274* .8579*
Llama 8B-1B
SLM w/o 7.04 6.55 33.20 18.20 28.55 31.75 14.92 19.78 20.81 10.21 17.30 .6535 .6355 .5900 .6646
CoSteer 7.29 7.73* 41.28* 24.97* 31.19* 31.68 13.68 24.57* 26.11 12.00 23.69* .8911* .8582* .7812* .8721*
@@
Table 7: Paired T-test results. Bold entries indicate that CoSteer outperforms the three baselines. An asterisk (*) denotes that this improvement is statistically significant (paired t-test, $p<0.05$). Gray values represent the privacy-violating near-upper-bound performance, yet underlined CoSteer values surpasses these incompatible references.
@@
To formally validate the robustness of our findings, we performed paired t-tests comparing CoSteer against each of the three key baselines: LLM w/o, SLM w/, and SLM w/o. The tests were conducted on the full set of evaluation samples, comparing the performance scores on a per-sample basis to determine if the observed improvements were due to more than random chance. We used a standard significance level of $\alpha=0.05$. The results confirmed that the performance gains of CoSteer over all three baselines are statistically significant. In Table 7, these significant improvements are denoted with an asterisk (*).
@@ B.4 Detailed Analysis on Experiment Results
Our framework demonstrates distinct performance patterns across task complexity levels. For context-intensive generation tasks (Cogenesis and Longlamp) requiring deep integration of extended user profiles, histories, and multiple writing examples, model pairs with larger size demonstrate superior contextual reasoning capabilities. Qwen 32B-7B shows significant gains across three datasets compared to compact models (Qwen 7B-1.5B/ Llama 8B-1B). In contrast, the Llama8B-1B configuration occasionally underperforms baselines by failing to distill essential personalization signals from sparse contexts. Conversely, in preference alignment tasks where personal context reduces to concise and explicit textual instructions, compact models perform comparably to their larger counterparts. This stems from the larger models’ over-alignment when processing simplistic personalization signals and effectively overfitting to sparse preference indicators.
@@
B.5 Dataset Descrition
@@ B.7 Prompts for LLM-as-a-Judge
We use exactly the same prompt as Cogenesis (Zhang et al., 2024b) to have the gpt-4o evaluate the generated content. The specific prompts are shown in Figure 6 and Figure 7.
@@ B.8 Baseline Implementation Details
Context Steering (CoS) (He et al., 2025) and Linear Alignment (LA) (Gao et al., 2024) are conceptually equivalent methods. They steer the generation of a base model by modifying its output log-probabilities. The final log-probability for a token $x_{i}$ is calculated by adding a scaled difference term to the base model’s prediction, which is conditioned on no context ($\emptyset$). The formulation is as follows:
@@
This method (Liu et al., 2024) enhances a large base model by incorporating signals from a smaller, specialized proxy model. The final probability distribution is calculated by adjusting the logits of the large model. We adapt the original formula to better reflect the roles of the different models in our setup:
@@ B.9 Parameter Sensitivity Analysis
Figure 2: Effect of different $\alpha$ , $\beta$ , $\eta$ , $\lambda$ and iteration time $T$ on the Abstract Generation dataset using Qwen 7B-1.5B. The evaluation metric is ROUGE-L.
@@ B.10 Robostness to Noise
To supplement the discussion in Section 5.1, this section provides the detailed experimental setup and result analysis for our robustness tests.
@@ B.10.2 Robustness to Purely Irrelevant (Adversarial) Noise
To push the limits of our framework, we designed a more adversarial scenario to test a potential failure mode: providing the model with context that is completely irrelevant to the task. For example, when generating an Abstract, we supplied context examples drawn from the Review task. This setup evaluates whether the model can ignore distracting, irrelevant information or if it gets catastrophically misled. We use Qwen7B-1.5B pair to conduct the experiment.
@@
Table 10 shows the outcomes of this stress test. As expected, the performance of all models degrades compared to using relevant context. However, the key observation is in the *manner* of degradation. CoSteer (e.g., 39.59 R-1 on Abstract) experiences a controlled performance drop but still comfortably outperforms the specialized SLM w/ (34.91 R-1). It does not collapse. This graceful degradation suggests that the strong inductive bias from the frozen base LLM acts as a safeguard, preventing the steering mechanism from being entirely derailed by nonsensical context. The model learns to rely more on its pretrained knowledge when the provided context is useless, showcasing a desirable level of robustness.
@@ B.11 Results and Implementation of Cross-Architecture Collaboration
Table 11: Performance of CoSteer in a cross-architecture setting, using Llama-3.1-8B as the LLM and Qwen2.5-1.5B as the SLM. We compare two vocabulary-agnostic strategies: vocabulary mapping (CoSteer_map) and a more universal byte-level fusion (CoSteer_byte). The results show that both methods enable effective collaboration, validating our framework’s generalization capability across different model families.
@@
The experiments were conducted using Llama-3.1-8B as the LLM and Qwen2.5-1.5B as the SLM. Results are shown in Table 11. Below we details the two strategies we implemented to facilitate collaboration between models with different architectures and tokenizers, as discussed in Section 5.2.
@@ B.11.2 Byte-level Fusion (CoSteer_byte)
To overcome the limitations of vocabulary mapping, CoSteer_byte operates entirely in a shared byte space. Instead of aligning token IDs, it projects each model’s token-level logit distribution into a common 257-dimensional space (Hayase et al., 2025): 256 dimensions for possible next bytes (0–255) and one special “commit” dimension (byte 256) indicating that the current token is complete. This enables direct byte-level probabilistic fusion of LLM and SLM outputs.
@@ B.12.4 Efficiency and Overhead Analysis
To understand the sources of the observed latency, we analyze the design of our variants and the nature of the overhead. We proposed LightCoSteer and AdaCoSteer to systematically explore the trade-offs between generation quality and efficiency.
  • LightCoSteer simplifies the framework by setting the number of optimization iterations $T=1$. This eliminates the iterative process, effectively reducing the mechanism to a single-step logit adjustment. This variant serves as an ablation study on the Follow-the-Regularized-Leader (FTRL) optimization itself.
  • AdaCoSteer is motivated by our observation that the LLM’s predictions tend to converge with the SLM’s guidance as generation progresses. It implements an adaptive termination strategy (Song et al., 2025): the steering process is halted once the LLM’s output confidence (the probability of the token selected via greedy search) remains above a predefined threshold for $c$ consecutive steps. After this condition is met, generation continues using only the remote LLM, which eliminates all further communication overhead.
This design allows us to qualitatively break down the latency of a single token generation into three components: the LLM inference latency $\mathcal{L}(n)$, the data transmission duration $\mathcal{T}(n)$, and the per-token optimization cost $\mathcal{C}(n)$. The empirical results lead to two key conclusions:
  1. 1.
    Optimization cost is minimal: The small speed difference between the iterative CoSteer and the single-step LightCoSteer confirms that the computational cost of the optimization, $\mathcal{C}(n)$, is negligible. This is because our closed-form solution relies on highly efficient, basic tensor arithmetic.
  2. 2.
    Transmission is the primary bottleneck: The significant latency gap between vanilla LLM generation and even the fast LightCoSteer variant highlights that the data transmission time, $\mathcal{T}(n)$, is the dominant source of overhead in our framework.
In conclusion, our analysis confirms that the primary overhead is from communication, not the optimization algorithm itself. Adaptive strategies like AdaCoSteer can effectively mitigate this bottleneck, offering users a flexible way to balance personalized generation with operational efficiency.
@@
Appendix C Broader Impact and Future Work
Broader Impact
The primary contribution of our work lies in addressing the critical tension between the advancement of large-scale AI and the fundamental right to personal privacy. By enabling powerful, cloud-based LLMs to be personalized using sensitive data that never leaves the user’s device, the CoSteer framework offers a practical and scalable solution for privacy-preserving AI. This aligns directly with societal demands for data sovereignty and user-centric control.
Furthermore, our approach represents a significant step forward in the field of AI alignment. Instead of relying on a single, universal alignment target, CoSteer facilitates a more dynamic and democratic form of alignment where model behavior is adapted in real-time to the unique context, preferences, and style of the individual. This fosters a safer and more beneficial human-AI interaction paradigm, reducing the risk of generating generic or contextually inappropriate content. By decentralizing the personalization process, our work contributes to a more equitable technological ecosystem where users can benefit from state-of-the-art AI without compromising their data.
Future Work
The CoSteer framework establishes a versatile foundation for a new class of collaborative, privacy-aware generative models. Looking ahead, we plan to explore the broader potential of this paradigm.
Appendix D LLM Usage
The authors acknowledge the use of large language models (LLMs) during the composition of this paper. The models’ role was confined to that of an assistive tool for language refinement, including grammar correction, stylistic improvements, and punctuation adjustments. It is important to state that the LLMs did not contribute to the generation of any original scientific concepts, experimental methodologies, or the analysis and interpretation of results. The intellectual contributions are entirely those of the human authors.

### v3

@@ A.2 Proof
Appendix B Experiment Details
@@ B.1 Hyperparameters and Settings
To ensure the results are easily reproducible, we set do_sample=False for all methods. For the Qwen3 series, we set enable_thinking=False. Hyperparameters of CoSteer are listed in Table 4
@@ B.3 Statistical Significance Testing
Qwen 7B-1.5B SLM w/o 6.63±1.70 6.21±1.88 36.48±7.50 17.74±3.85 27.57±5.44 20.40±8.88 10.39±3.71 10.39±6.07 25.21±9.55 11.09±3.24 17.51±8.08 .5441±.13 .5154±.11 .6104±.09 .5632±.06
SLM w/ 7.81±1.24 7.63±1.43 39.75±9.83 22.03±9.83 27.76±10.03 23.08±9.95 12.40±7.15 12.94±8.90 22.89±9.19 11.46±6.30 17.12±8.45 .6214±.10 .6998±.08 .7102±.13 .6594±.05
LLM w/o 8.00±1.18 7.63±1.43 39.81±6.86 20.53±4.32 25.56±7.11 30.15±6.64 14.04±2.86 17.71±6.54 27.64±8.70 11.93±2.48 21.49±5.51 .7058±.10 .6931±.08 .7039±.13 .7183±.07
CoSteer 8.44*±1.34 8.50*±1.20 42.98*±8.94 23.61*±7.09 28.20*±8.88 32.72*±9.67 15.92*±9.02 20.36*±10.09 25.93±10.13 12.38*±9.78 22.84*±10.41 .7915*±.06 .7608*±.07 .7404*±.08 .7885*±.14
LLM w/ 8.62±0.63 8.61±0.58 44.50±9.77 24.63±9.14 31.15±9.34 33.83±5.44 15.55±2.41 22.42±4.76 28.82±11.76 13.77±9.37 23.44±8.29 .8884±.07 .8076±.05 .7693±.07 .8414±.16
Qwen 32B-7B SLM w/o 8.00±1.18 7.63±1.43 39.81±6.86 20.53±4.32 25.56±7.11 30.15±6.64 14.04±2.86 17.71±6.54 27.64±8.70 11.93±2.48 21.49±5.51 .7058±.09 .6931±.09 .7039±.08 .7183±.15
SLM w/ 8.62±0.63 8.61±0.58 44.50±9.77 24.63±9.14 31.15±9.34 33.83±5.44 15.55±2.41 22.42±4.76 28.82±11.76 13.77±9.37 23.44±8.29 .8884±.05 .8076±.06 .7693±.12 .8414±.05
LLM w/o 8.12±1.01 7.87±1.16 40.66±6.79 21.02±4.26 26.61±6.83 32.21±6.32 14.44±2.51 19.61±6.44 28.82±8.53 12.20±2.68 21.16±6.11 .7020±.14 .6873±.06 .7225±.10 .7146±.13
CoSteer 8.78*±0.54 8.64*±0.79 45.41*±9.79 26.04*±9.99 33.52*±9.64 34.88*±8.80 15.89*±4.72 26.51*±7.23 30.10*±12.05 14.52*±10.76 24.20*±11.26 .8589±.06 .8532*±.08 .7274±.06 .8579*±.04
LLM w/ 8.83±0.48 8.76±0.51 43.33±8.16 23.47±6.40 30.10±7.36 34.65±5.63 15.74±3.83 22.77±5.48 30.73±10.83 14.20±8.45 24.25±8.67 .9017±.03 .8193±.07 .7912±.12 .8538±.06
Llama 8B-1B SLM w/o 7.04±1.65 6.55±1.81 33.20±8.60 18.20±3.78 28.55±4.84 31.75±5.95 14.92±2.37 19.78±5.53 20.81±10.10 10.21±3.22 17.30±8.29 .6535±.12 .6355±.11 .5900±.10 .6646±.12
SLM w/ 7.69±1.40 7.52±1.39 39.81±9.63 21.53±4.95 30.11±6.45 32.36±6.59 15.02±2.70 22.06±4.80 20.17±10.43 10.58±5.99 18.64±8.61 .7981±.11 .7908±.08 .7037±.13 .8065±.10
LLM w/o 7.69±1.38 7.13±1.62 39.33±8.11 20.69±3.90 29.41±5.32 34.58±5.19 15.32±2.43 22.10±5.54 26.93±10.11 12.35±3.40 21.82±6.47 .7038±.09 .6878±.06 .6765±.09 .7116±.09
CoSteer 7.29±1.42 7.73*±1.03 41.28*±8.36 24.97*±8.10 31.19*±6.42 31.68±6.33 13.68±4.63 24.57*±6.93 26.11±11.18 12.00±6.74 23.69*±8.13 .8911*±.05 .8582*±.08 .7812*±.12 .8721*±.10
LLM w/ 8.61±0.66 8.44±0.80 43.91±9.69 23.93±6.44 32.01±7.97 36.39±5.19 15.95±2.50 23.56±4.80 30.54±10.43 14.02±6.15 23.81±7.39 .8869±.06 .8298±.05 .7909±.07 .8551±.06
Qwen 8B-0.6B SLM w/o 6.84±1.70 6.64±1.80 37.13±7.17 20.41±4.90 21.00±6.91 24.41±7.88 12.83±3.69 12.39±5.68 24.19±9.46 11.88±3.88 15.74±6.38 .5710±.13 .5359±.04 .5981±.10 .5937±.11
SLM w/ 7.85±1.49 7.79±1.42 42.48±8.39 23.25±6.21 26.58±8.29 25.04±10.20 13.36±7.38 13.02±8.71 26.17±11.92 13.46±9.01 18.40±10.13 .6386±.12 .6432±.14 .6000±.14 .6436±.12
LLM w/o 8.27±1.05 7.99±1.25 40.76±7.06 21.23±4.37 25.77±6.75 30.89±6.69 14.47±2.85 16.98±5.79 28.47±9.30 12.85±2.94 21.49±6.65 .7597±.15 .7530±.14 .6929±.15 .7611±.07
CoSteer 8.62*±0.90 8.65*±1.12 41.11±8.51 22.36±6.21 25.86±7.98 32.24*±7.22 14.84*±3.02 18.47*±6.15 29.03*±11.07 13.61*±7.06 22.98*±8.47 .7619*±.06 .7573*±.08 .6913±.05 .7648*±.11
LLM w/ 8.96±0.46 8.96±0.36 43.99±8.46 23.69±6.42 29.48±7.76 35.42±5.45 15.71±2.55 21.59±4.96 31.48±13.00 15.34±11.50 24.78±11.46 .8895±.06 .8031±.12 .7713±.09 .8552±.05
Qwen 32B-0.6B SLM w/o 6.84±1.70 6.64±1.80 37.13±7.17 20.41±4.90 21.00±6.91 24.41±7.88 12.83±3.69 12.39±5.68 24.19±9.46 11.88±3.88 15.74±6.38 .5710±.06 .5359±.10 .5981±.12 .5937±.06
SLM w/ 7.85±1.49 7.79±1.42 42.48±8.39 23.25±6.21 26.58±8.29 25.04±10.20 13.36±7.38 13.02±8.71 26.17±11.92 13.46±9.01 18.40±10.13 .6386±.08 .6432±.11 .6000±.10 .6436±.12
LLM w/o 8.31±0.92 8.22±1.12 41.05±6.95 21.88±4.54 26.23±7.01 30.23±6.94 14.08±3.03 16.41±5.75 29.01±10.00 12.51±3.06 21.44±6.23 .7861±.10 .7850±.08 .6934±.10 .7783±.06
CoSteer 8.52*±0.89 8.56*±0.87 43.74*±9.51 23.69*±8.47 30.97*±9.74 30.89*±9.37 14.14±4.89 18.87*±7.62 26.75±12.75 13.17±10.14 21.32±8.47 .7858±.11 .7883±.09 .6873±.14 .7790±.07
LLM w/ 9.09±0.44 9.03±0.29 44.46±8.08 23.95±5.47 30.15±6.98 35.12±6.20 15.60±2.70 20.65±5.45 32.63±12.44 15.37±10.48 25.07±10.63 .9095±.05 .8095±.08 .8040±.13 .8654±.06
@@
Table 6: Main Results with Mean and Standard Deviation. Bold entries indicate that CoSteer outperforms the three baselines (SLM w/o, SLM w/, LLM w/o). An asterisk (*) denotes that this improvement is statistically significant (paired t-test, $p<0.05$). Gray values represent the privacy-violating near-upper-bound performance, yet underlined CoSteer values surpass these incompatible references.
@@
To formally validate the robustness of our findings, we performed paired t-tests comparing CoSteer against each of the three key baselines: LLM w/o, SLM w/, and SLM w/o. The tests were conducted on the full set of evaluation samples, comparing the performance scores on a per-sample basis to determine if the observed improvements were due to more than random chance. We used a standard significance level of $\alpha=0.05$. The results confirmed that the performance gains of CoSteer over all three baselines are statistically significant. In Table 6, asterisks (*) denote these significant improvements.
@@ B.4 Detailed Analysis on Experiment Results
Our framework demonstrates distinct performance patterns across task complexity levels. For context-intensive generation tasks (Cogenesis and LongLaMP) requiring deep integration of extended user profiles, histories, and multiple writing examples, larger model pairs demonstrate superior contextual reasoning capabilities. Qwen 32B-7B shows significant gains across three datasets compared to compact models (Qwen 7B-1.5B/ Llama 8B-1B). In contrast, the Llama8B-1B configuration occasionally underperforms baselines by failing to distill essential personalization signals from sparse contexts. Conversely, in preference alignment tasks where personal context reduces to concise and explicit textual instructions, compact models perform comparably to larger counterparts. This stems from the larger models’ over-alignment when processing simplistic personalization signals and effectively overfitting to sparse preference indicators.
@@
B.5 Dataset Description
To ensure a comprehensive evaluation relevant to Cloud-LLM serving scenarios, we selected datasets covering a broad spectrum of personal AI agent applications:
  • Daily Assistance (Cogenesis): This benchmark simulates daily tasks such as drafting emails and notifications based on user profiles and history.
  • Complex Content Creation (LongLaMP): This focuses on reasoning-intensive, long-form generation tasks including academic abstracts, product reviews, and blog posts.These tasks typically require the superior generation capabilities of Cloud LLMs to ensure high quality.
  • General Preference Alignment: We use four datasets (e.g., HelpSteer, TruthfulQA) to evaluate general instruction following based on specific user constraints.
@@ B.7 Prompts for LLM-as-a-Judge
We use exactly the same prompt as Cogenesis (cogenesis) to have the gpt-4o evaluate the generated content. The specific prompts are shown in Figure 7 and Figure 8.
@@ B.8 Baseline Implementation Details
Context Steering (CoS) (he2025contextsteeringcontrollablepersonalization) and Linear Alignment (LA) (gao2024linearalignmentclosedformsolution) are conceptually equivalent methods. They steer the generation of a base model by modifying its output log-probabilities. The final log-probability for a token $x_{i}$ is calculated by adding a scaled difference term to the base model’s prediction, which is conditioned on no context ($\emptyset$). The formulation is as follows:
@@
This method (liu2024tuninglanguagemodelsproxy) enhances a large base model by incorporating signals from a smaller, specialized proxy model. The final probability distribution is calculated by adjusting the logits of the large model. We adapt the original formula to better reflect the roles of the different models in our setup:
@@
The comparison results are shown in Table 8.
Models Method Abstract Review Writing
R-1 R-L MET R-1 R-L MET R-1 R-L MET
Qwen 7B-1.5B LA/CS 41.49 25.57 29.96 26.60 13.68 17.59 24.09 12.32 19.94
SFT 40.71 21.85 27.00 31.55 12.91 20.23 25.53 10.80 19.35
Proxy tuning 40.76 21.37 25.95 29.54 14.16 16.79 27.13 12.11 21.38
Costeer 42.98 23.61 28.20 32.72 15.92 20.36 25.93 12.38 22.84
Qwen 32B-7B LA/CS 44.28 25.85 32.98 34.14 15.46 26.38 29.22 14.03 23.66
SFT 43.59 25.48 32.43 31.94 11.84 21.00 28.83 10.63 21.73
Proxy tuning 44.61 25.12 31.97 32.03 14.80 19.01 28.75 12.70 21.49
Costeer 45.41 26.04 33.52 34.88 15.89 26.51 30.10 14.52 24.20
Llama 8B-1B LA/CS 37.24 20.70 30.78 27.35 11.93 23.54 19.28 9.30 18.26
SFT 40.62 26.67 25.64 29.93 10.17 21.51 26.11 8.99 22.56
Proxy tuning 39.55 20.78 29.60 30.50 12.29 21.81 26.00 12.29 21.65
Costeer 41.28 24.97 31.19 31.68 13.68 24.57 26.11 12.00 23.69
Qwen 8B-0.6B LA/CS 31.42 16.78 22.10 28.37 12.92 15.94 25.31 11.92 18.96
SFT 38.38 21.88 26.03 30.53 14.75 19.91 25.20 12.82 18.83
Proxy tuning 40.60 21.14 25.68 30.97 14.48 16.94 28.84 12.89 21.64
Costeer 41.11 22.36 25.86 32.24 14.84 18.47 29.03 13.61 22.98
Qwen 32B-0.6B LA/CS 31.42 16.78 22.10 28.37 12.92 15.94 25.31 11.92 18.96
SFT 38.38 21.88 26.03 30.53 14.75 18.91 25.20 12.82 18.83
Proxy tuning 41.17 21.88 26.40 30.20 14.14 16.27 26.08 12.61 21.67
Costeer 43.74 23.69 30.97 30.89 14.14 18.87 26.75 13.17 21.32
Table 8: Performance on LongLamp benchmark across 5 model pairs. Best results are marked in bold and second-best are underlined.
@@ B.9 Parameter Sensitivity Analysis
Figure 2: Effect of $\alpha$ , $\beta$ , $\eta$ , $\lambda$ and iteration step $T$ on the Abstract Generation dataset using Qwen 7B-1.5B. Metric: ROUGE-L.
@@ B.10 Robostness to Noise
As a supplement to Section 5.1, we detail the experimental setup and results for our robustness tests.
@@ B.10.2 Robustness to Purely Irrelevant (Adversarial) Noise
To push the limits of our framework, we designed a more adversarial scenario to test a potential failure mode: providing the model with context that is completely irrelevant to the task. For example, when generating an Abstract, we supplied context examples drawn from the Review task. This setup evaluates whether the model can ignore distracting information or if it gets catastrophically misled. We use Qwen7B-1.5B pair to conduct the experiment.
@@
Table 10 shows the outcomes of this stress test. As expected, the performance of all models degrades compared to using relevant context. However, the key observation is in the *manner* of degradation. CoSteer (e.g., 39.59 R-1 on Abstract) experiences a controlled performance drop but still comfortably outperforms the specialized SLM w/ (34.91 R-1). It does not collapse. This graceful degradation suggests that the strong inductive bias from the frozen base LLM acts as a safeguard, preventing the steering mechanism from being entirely derailed by nonsensical context. The model learns to rely more on its pretrained knowledge when context is useless, showcasing a desirable level of robustness.
@@ B.11 Results and Implementation of Cross-Architecture Collaboration
Table 11: Performance of CoSteer in a cross-architecture setting, using Llama-3.1-8B as the LLM and Qwen2.5-1.5B as the SLM. We compare two vocabulary-agnostic strategies: vocabulary mapping (CoSteer_map) and a more universal byte-level fusion (CoSteer_byte). Both methods enable effective collaboration, validating our framework’s generalization capability across model families.
@@
The experiments were conducted using Llama-3.1-8B as the LLM and Qwen2.5-1.5B as the SLM. Results are shown in Table 11. Below we detail the two strategies we implemented to facilitate collaboration between models with different architectures and tokenizers, as discussed in Section 5.2.
@@ B.11.2 Byte-level Fusion (CoSteer_byte)
To overcome the limitations of vocabulary mapping, CoSteer_byte operates entirely in a shared byte space. Instead of aligning token IDs, it projects each model’s token-level logit distribution into a common 257-dimensional space (hayase2025byte-sampling): 256 dimensions for possible next bytes (0–255) and one special “commit” dimension (byte 256) indicating that the current token is complete. This enables direct byte-level probabilistic fusion of LLM and SLM outputs.
@@ B.12.4 Efficiency and Overhead Analysis
Figure 3: Detailed breakdown of wall-clock time per token. The system employs an asynchronous pipelining strategy: The Local SLM inference (Stream 2) is executed simultaneously with the Network/Cloud path (Stream 1). Since the local inference time ($\approx 30$ms) is typically shorter than the network round-trip ($\approx 40$ms) plus cloud inference ($\approx 40$ms), the local computational burden is effectively masked and does not affect the critical path latency. The primary bottlenecks are thus Network Transmission (b & e) and the local FTRL Optimization (d).
@@
To clearly identify the sources of overhead and justify our design choices, we provide a detailed breakdown of the wall-clock time per token in Figure 3. This visualization highlights the asynchronous pipelining workflow inherent to CoSteer.
As illustrated in Figure 3, the generation process is split into two parallel streams once a token is sampled:
  • Stream 1 (Critical Path): Involves uploading the token, Cloud LLM inference, and downloading logits. This path is dominated by network transmission ($\mathcal{T}(n)\approx 40$ms) and cloud inference ($\mathcal{L}(n)\approx 40$ms).
  • Stream 2 (Masked Path): Simultaneously, the local SLM performs batched inference. Crucially, because the local NPU inference time ($\approx 30$ms) is masked by the longer duration of Stream 1, it does not impose a penalty on the total latency.
Based on this breakdown, we proposed LightCoSteer and AdaCoSteer to systematically address the actual unmasked bottlenecks on the critical path:
  • LightCoSteer targets the Optimization Bottleneck (d). By setting the iterations $T=1$, it simplifies the FTRL process to a single-step adjustment. This variant confirms that while the iterative optimization ($\approx 25$ms) adds overhead, it is manageable.
  • AdaCoSteer targets the Transmission Bottleneck (b & e and together with d since no additional fusion are needed). Motivated by the observation that LLM and SLM predictions converge over time, it employs an adaptive termination strategy (song-etal-2025-well-begun). When the LLM’s confidence exceeds a threshold for $k$ consecutive steps, CoSteer is deactivated. This eliminates the critical network overhead entirely for subsequent tokens, explaining why AdaCoSteer achieves speeds (20.65 tok/s) comparable to vanilla generation.
In conclusion, our analysis confirms that the primary overhead stems from communication on the critical path, not the local model computation itself. Adaptive strategies like AdaCoSteer effectively mitigate this bottleneck, offering a flexible balance between personalized generation and operational efficiency.